OpenStudy (hesan):

A canister of gas was connected to a gas syringe and the wall open slightly to allow some of the gas into the syringe. It was found that 0.200 gm of the gas took up the volume of 96 cubic cm at a temp of 20 degree Celsius and a pressure of 1.02x10^5. Calculate a relative molecular mass of a gas mixture.

OpenStudy (hesan):

@ghazi

OpenStudy (hesan):

@kugler97

OpenStudy (aaronq):

rearrange ideal gas law.. PV=nRT sub in: n=m/M P=mRT/VM M=mRT/PV sub in: m/V=density M=dRT/P you're given mass and volume.. d=mass/volume you're given temperature and pressure plug and chug

OpenStudy (hesan):

the pressure is 1.02x10^5 pascal.. do i have to change it into atm? On changing it to atm, i m getting the answer 0.0497 kg . Is this correct?

OpenStudy (aaronq):

you have to change the temp to Kelvin and the pressure to atm the ideal gas law (or van der waals equation) always uses these units

OpenStudy (hesan):

ok let me calculate the answer now.. will u check whether it is correct or not?

OpenStudy (hesan):

and what about cubic cm? do i have to change it into cubic m or not?

OpenStudy (aaronq):

yeah i forgot you also have to change volume to L, which is cubic meters. sure, post your steps and il check it

OpenStudy (hesan):

P= 1.00674 atm V =0.096 cubic dm m = 0.200 gm R = 0.0821 atm dm^3 mol^-1 K^-1 T = 293 K Now, PV=m/M RT M= (mRT)/ PV M= (0.200 x 0.0821 x 293) / (1.00674 x 0.096) M= 49.59 g/mol or M= 0.0495 kg/mol

OpenStudy (aaronq):

i got 49.779 which can be rounded up to 49.8 g/mol

OpenStudy (aaronq):

but regardless, you got the steps right.